A car moving along a straight road at a velocity of 30ms-¹ reduces its velocity at a constant rate until it stop after 5 s. What is the acceleration of the car? 


1
Expert's answer
2021-02-15T17:37:58-0500
"a=\dfrac{v-v_0}{t}=\dfrac{0-30\ \dfrac{m}{s}}{5\ s}=-6.0\ \dfrac{m}{s^2}."

The sign minus means that the car is decelerates.